OpenStudy (anonymous):

Find the inverse function in slope-intercept form (mx+b): f(x)=5/3x+5

OpenStudy (anonymous):

is this \[f(x)=\frac{5}{3}x+5\]?

OpenStudy (anonymous):

Yes, I need helpp.

OpenStudy (anonymous):

ok we can do this it is just like solving \[\frac{5}{3}x+5=7\] more or less

OpenStudy (anonymous):

but this is more general start with \[y=\frac{5}{3}x+7\] then switch \(x\) and \(y\) to write \[x=\frac{5}{3}y+7\] and then solve for \(y\) it will take two steps

OpenStudy (anonymous):

How would i start solving it like that ?

OpenStudy (anonymous):

\[x=\frac{5}{3}y+7\] subtract \(7\) from both sides to get \[x-7=\frac{5}{3}y\] multiply by 3 and get \[3x-21=5y\] divide by \(5\) to get \[\frac{3x-21}{5}=y\]

OpenStudy (anonymous):

and that is your inverse: \[f^{-1}(x)=\frac{3x-21}{5}\]

OpenStudy (anonymous):

oh if you want it in "slope intercept" form write it as \[f^{-1}(x)=\frac{3}{5}x-\frac{21}{5}\]

OpenStudy (anonymous):

The correct answer was 3/5x-3

OpenStudy (anonymous):

did i make a mistake somewhere?

OpenStudy (anonymous):

oh my fault there is no seven in the problem it is \[x=\frac{5}{3}y+5\]

OpenStudy (anonymous):

ok same idea \[x-5=\frac{5}{3}y\] \[3x-15=5y\] \[\frac{3x-15}{5}=y\] or \[y=\frac{3}{5}x-3\]